OpenStudy (anonymous):

okay, so i know the quotient rule and chain rule but im confused on when applying it. the question is find the derivative of the function. y=(x^2+1 / x+2-1)^3 I used chain rule and got 3(x^2+1 / x+2-1)^2 times (2x/2x) I am then told to use quotient rule. can someone help me step by step please ?!

OpenStudy (phi):

is this what is in the parens: \[ \frac{x^2+1}{x+2} -1 \]?

OpenStudy (anonymous):

ah im sorry ! the denominator is x^2-1

OpenStudy (phi):

so the problem is \[ \frac{x^2+1}{x^2-1} \]? the rule is bottom * derivative top - top * derivative bottom all divided by the bottom squared. the top of the derivative is \[ (x^2 - 1) (2x) - (x^2+1)(2x) \\ 2x^3 -2x - 2x^3 - 2x \\ - 4x \] so the answer is \[ \frac{ -4x}{(x^2-1)^2} \]